1.Pipe A can fill a tank in 30 minutes. Pipe B can feel the same tank in 20 minutes. How long will it take to fill the tank using both pipes at the same time?
NOTE:( can you please show the work or somehow explain them so i can understand how to do it..thank you)

Answer by stanbon(75887)

Pipe A can fill a tank in 30 minutes. Pipe B can feel the same tank in 20 minutes. How long will it take to fill the tank using both pipes at the same time?
-----------------
A rate: 1/30 tank/min
B rate: 1/20 tank/min
-------------------------
Together rate: 1/x tank/min
==================================
Equation:
rate + rate = together rate
1/30 + 1/20 = 1/x
--------
Multiply thru by 60x and solve for "x":
2x + 3x = 60
5x = 60
x = 12 minutes (time to fill the tank using both tanks)
============================================================
Cheers,
Stan H.